# GRAVITY for MATISSE

### v1

@@ 4.3 Reduced N-band coherent flux bias
Due to the very short exposures and the large thermal background noise, the MATISSE N-band fringes must be coherently integrated in post-processing. When MATISSE operates independently, this coherent integration is implemented in the pipeline (Lopez et al., 2022, section 3.3) and relies on an estimation of the phase from the same data being coherently integrated. This operation generates a bias at low flux levels, as illustrated in Fig. 7. When operating with the fringe tracker, the fringe phase is measured from the higher SNR K-band data, extrapolated to the N band following the method of section 2.8. In these conditions, the bias is eliminated. The use of GRA4MAT is therefore always recommended below with the ATs and below with the UTs, as it allows reaching the respective fundamental (photon, thermal, readout) noise limits of and .
4.4 Stabilized LM-bands transfer function
The use of GRA4MAT makes the MATISSE LM-bands transfer function much less sensitive to seeing and coherence time conditions, as illustrated in Fig. 8. The fringe tracker maintains the LM-bands instrumental contrast at a level of 0.9 down to a very low coherence time of , whereas the contrast of MATISSE standalone starts decreasing below a coherence time value of about . Seeing values above about show a decrease of the transfer function stability in MATISSE standalone and with GRA4MAT but more marginally.
Fig. 8 should however not be interpreted as a demonstration that GRA4MAT can be used in poor seeing conditions. The contrasts shown in the figure were selected for MATISSE DITs without fringe jumps. We recall (see section 3) that the fringe tracker performance depends on the delivery of a stable Strehl by the adaptive optics.
From the dataset presented in Fig. 8, we can estimate the raw visibility accuracy as the variability of the contrast in changing atmospheric conditions. For the 50% average atmospheric conditions, seeing ¡ and coherence time larger than , the standard deviation of the contrast is estimated at ±0.08 for MATISSE, and ±0.02 for GRA4MAT. The MATISSE value matches the estimation given in the first light paper (Lopez et al., 2022, Table 5), whereas the GRA4MAT value bests to MATISSE standalone performance in good atmospheric conditions. Additional information on the raw contrast performance is given in Table 4.
@@ 4.5 Visibility measurements
Absolute visibility measurements with MATISSE require accurate photometry and hence chopping at frequencies higher than for targets fainter than at and at 4 µm on ATs (respectively and with UTs). This cannot be improved by GRA4MAT. Only the DITs for the LOW spectral resolution of MATISSE are compatible with this chopping frequency. However, absolute visibility measurements in the continuum are mandatory only in the same LOW spectral resolution. At higher spectral resolutions, MATISSE users are mostly looking for differential measures between spectral features that are not affected by broad-band photometric errors. Hence, absolute visibility measurements with GRA4MAT should be restricted to LOW resolution, while higher resolutions with long DITs and without chopping should be used only for coherence flux and differential measurements. Simultaneous chopping and fringe tracking was developed for GRA4MAT in this context (see section 2.6).
@@ 5 Outlook
The performance of GRAVITY for MATISSE presented here corresponds to what was achieved at the end of the project in mid-2023, and justify the recommendation to always use GRA4MAT over MATISSE standalone whenever possible.
@@ 5.1 Science highlight: a first exoplanet
Soon after the first exoplanet discovery (Mayor & Queloz, 1995), their observation by differential phase interferometry333Other interferometric techniques like indirect detection by wide angle astrometry (Shao & Colavita, 1992; Colavita et al., 1998, 1999; Delplancke, 2008; Woillez et al., 2010) and direct detection by nulling (Serabyn et al., 2012; Absil et al., 2006) were also proposed and explored, but not developed herein. was considered (Akeson & Swain, 1999; Petrov et al., 2001). Detection attempts based on the differential phase technique were carried out at VLTI, with the the near-infrared AMBER instrument (Millour et al., 2006; Vannier et al., 2006). In the mid-infrared specifically, a first attempt was made with simultaneous MIDI and AMBER observations (Matter et al., 2010). The goals were to evaluate and calibrate the chromatic effects of air, strongly impacting the mid-infrared domain, and to prepare the approach for future MATISSE observations in this field. In addition, it was realised at that time that parasitic light effects would introduce biases affecting the planet signal measurement (Matter et al., 2009); MATISSE was designed to prevent such parasitic effects. The scientific requirements and conceptual study of MATISSE were taking into account the exigent needs imposed by high contrast measurements on exoplanets (Lopez et al., 2014; Wolf et al., 2016).
@@
The high sensitivity potential of the off-axis fringe tracking with GRA4MAT was demonstrated on the planet Pictoris b with MATISSE during the commissioning of the GRA4MAT narrow off-axis mode, on the nights of November 8, 2022 and February 3, 2023 (Prog. ID 60.A-9257). These observations were motivated by an earlier observing Guaranteed Observing Time proposal (”Retrieving the Pictoris b spectrum in L band”, Prog. ID 110.23SJ), submitted to the ESO Observing Program Committee in September 2022, but not accepted because the mode had not yet been implemented.
We selected the host star Pictoris a as the fringe-tracking guide star, and alternatively observed the host star and the planet, at the offset predicted by whereistheplanet
@@
Fig. 9 shows the fringes obtained on the planet position where the contributions of the L= planet and of the star are nearly equivalent.
This corresponds to an attenuation factor of for this L= host star at ( at ).
A preliminary data reduction derived from the K-band GRAVITY approach of GRAVITY Collaboration et al. (2020), yields a first estimate of the planet spectrum.
To check consistency with a detection of Pictoris b, we use species
@@
We then compare our preliminary L-band spectrum to this best fit, extrapolated over the L-band.
@@ 5.2 Planned improvements
The performance of GRA4MAT will continue to improve due to imminent changes to the VLTI infrastructure for the UTs. GPAO, the new adaptive optics for VLTI/UT developed by the GRAVITY+ project (GRAVITY+ Collaboration et al., 2022a), will significantly enhance the behaviour of GRA4MAT on the large telescopes. Anticipated in 2024, the Natural Guide Star (NGS) mode will deliver significantly higher Strehl on brighter objects and make the instrument much less sensitive to degraded seeing conditions. We expect the fringe tracker operation to become fringe-jump-free on the UTs, just as it currently is on the ATs. When the Laser Guide Star (LGS) mode of GPAO becomes available, good Strehl conditions will be obtained on targets much fainter in the visible. The fringe tracking limiting magnitude mismatch between ATs and UTs will then be remedied. At a foreseen limiting magnitude of K=, and virtually no limit on the Adaptive Optics (AO) tip/tilt star at GAIA G=, GRA4MAT will be able to systematically observe red and faint targets like Young Stellar Objects and Active Galactic Nuclei.
We also expect that the higher and more stable Strehl delivered to the fringe tracker will make GRA4MAT compatible with chopping on the UTs. This capability will enable the high precision visibility measurements, in MEDIUM resolution and above, required for sensitive mid-infrared interferometric imaging with the larger telescopes.
As explained in section 2.3, GRA4MAT uses the main VLTI delay lines as common-path piston actuators, and suffers from a latency penalty (see Fig. 3). This penalty is about to be removed by using the Beam Compressor Differential Delay Lines recently installed by the GRAVITY+ project. These additional common-path delay lines, primarily developed as improved Differential Delay Lines for GRAVITY wide off-axis observations (GRAVITY+ Collaboration et al., 2022b), provide the same level of performance as the TTP internal actuators of the GRAVITY fringe tracker.
@@ 5.3 Possible future developments
As the time of writing this paper, with the exception of section 5.2, the GRA4MAT project at ESO is concluded. What follows are possible developments that should be considered.
@@
The GRA4MAT implementation could benefit from an additional fast piston actuator not in the common path but exclusive for MATISSE. With such a setup, the GRAVITY fringe tracker could send an open-loop feed-forward command, improving further the OPD rejection at low frequencies. This control strategy was prototyped (Lane & Colavita, 2003) on the Palomar Testbed Interferometer (Colavita et al., 1999) and operated with the Keck Nuller (Colavita et al., 2010) on the Keck Interferometer (Colavita et al., 2013). The benefit for MATISSE, a higher and more stable fringe contrast, might not be as strong as for a high-precision nulling instrument like NOTT (Defrère et al., 2018).
GRA4MAT would also benefit from further developing the MATISSE pipeline. Coherent integrations in post-processing over durations well beyond the DITs are already possible in the N band. In the LM bands however, longer than coherent integrations must be implemented, especially for observations involving the more sensitive UTs. The beta Pictoris b results shown in figure 9 actually demonstrate that integrations beyond the DITs are possible. Based on Fig. 6, the instantaneous coherent flux SNR for this target is whereas the contrast spectrum derived from the combination of many DITs reaches an SNR of . We hope that one day, MATISSE becomes capable of measurements on targets virtually invisible in individual LM-bands DITs.
We believe that adding a wide off-axis capability to MATISSE has a strong scientific potential. A MATISSE wide implementation would not require any new hardware developments. The dual-field building blocks are already present at VLTI and actively used by GRAVITY wide (GRAVITY+ Collaboration et al., 2022b). MATISSE and the infrared tip/tilt sensor IRIS can already receive the off-axis beams while keeping GRAVITY on axis. MATISSE wide would just be an extension of the GRA4MAT control software and the commissioning of a new mode. Considering that the atmospheric turbulence retains coherence at larger off-axis distances in the mid-infrared, MATISSE wide would drastically expand the MATISSE sky coverage. We would then be able to start observation campaigns of emblematic AGN and YSO science cases (see e.g. Boskri et al., 2021).
@@ 6 Conclusions
The GRAVITY fringe tracker is now used by both the GRAVITY and MATISSE instruments. As a result, MATISSE saved costs and benefited from the fringe tracking expertise developed by the GRAVITY team. Likewise, GRAVITY benefited from a single focus on its fringe tracker, avoiding dispersion and effort duplication. For example, MATISSE motivated the fringe jump control feature, which is also an added value for GRAVITY for high precision differential visibility across the K band. Therefore, selecting one shared fringe tracker federated the teams around one single project, and made it better for all. This illustrates the value of a mutualisation strategy.
Thanks to GRA4MAT, MATISSE now delivers its predicted performance. A summary of all the performance improvements is given in Table 4, as well as the performance improvements expected with the UTs, when GRAVITY+’s GPAO becomes available in late 2024.
A second wave of astrophysical MATISSE observations involving sensitivity, spectroscopy, accuracy and contrast is now happening; associated publications are on their way. The first narrow off-axis observations with MATISSE are opening a new field of mid-infrared extrasolar planet observations with VLTI (Houllé et al., 2024).
For the past 20 years, single-telescope instrumentation for high angular resolution and high contrast observations has been unimaginable without the support of an adaptive optics system (Wizinowich et al., 2000; Brandner et al., 2002). A similar revolution is now happening for optical long-baseline interferometry. This paradigm shift reveals itself clearly when comparing early interferometry reviews (Quirrenbach, 2001; Monnier, 2003) to more recent ones (Eisenhauer et al., 2023). The interferometric instruments, that used to focus on bright targets within the tight coherence limits of the atmosphere, cannot be conceived any more without fringe tracking support. Initiated by the GRAVITY instrument (GRAVITY Collaboration et al., 2017), GRAVITY for MATISSE has now completed this revolution for all modern VLTI instruments.
@@ Acknowledgements.
Based on observations collected at the European Southern Observatory under ESO commissioning programme 60.A-9257. This research has made use of the Jean-Marie Mariotti Center Aspro service 444Available at http://www.jmmc.fr/aspro. P. Berio, M. Houllé, S. Lacour and F. Millour acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-21-CE31-0017 (project ExoVLTI). J. Leftley and R. Petrov acknowledge the support of the french ANR under the grant AGN_MELBa ANR ANR-21-CE31-001. A. Meilland is supported by the ANR grant MASSIF (ANR-21-CE31-0018). J. Varga is funded from the Hungarian NKFIH OTKA project no. K-132406. J. Varga has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 716155 (SACCRED). The research of J. Varga is supported by NOVA, the Netherlands Research School for Astronomy. J. Varga acknowledges support from the Fizeau exchange visitors program. The research leading to these results has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement 101004719 (ORP).References

### v2

@@ 4.3 Reduced N-band coherent flux bias
Due to the very short exposures and the high thermal background noise, the MATISSE N-band fringes must be coherently integrated in post-processing. When MATISSE operates independently, this coherent integration is implemented in the pipeline (Lopez et al., 2022, section 3.3) and relies on an estimation of the phase from the same data being coherently integrated. This operation generates a bias at low flux levels, as illustrated in Fig. 7. When operating with the fringe tracker, the fringe phase is measured from the higher S/N K-band data, extrapolated to the N band following the method described in Sect. 2.8. In these conditions, the bias is eliminated. The use of GRA4MAT is therefore always recommended below with the ATs and below with the UTs, as it allows the respective fundamental (photon, thermal, readout) noise limits of and to be reached.
4.4 Stabilised LM-band transfer function
The use of GRA4MAT makes the MATISSE LM-band transfer function much less sensitive to seeing and coherence time conditions, as illustrated in Fig. 8. The fringe tracker maintains the LM-band instrumental contrast at a level of 0.9 down to a very low coherence time of , whereas the contrast of MATISSE standalone starts decreasing below a coherence time value of about . Seeing values above about show a decrease in the transfer function stability in MATISSE standalone and with GRA4MAT but more marginally.
Figure 8 should not be interpreted as a demonstration that GRA4MAT can be used in poor seeing conditions, however. The contrasts shown in the figure were selected for MATISSE DITs without fringe jumps. We recall (see section 3) that the fringe tracker performance depends on the delivery of a stable Strehl by the adaptive optics.
From the dataset presented in Fig. 8, we can estimate the raw visibility accuracy as the variability of the contrast in changing atmospheric conditions. For the 50% average atmospheric conditions, seeing ¡ and coherence time longer than , the standard deviation of the contrast is estimated at ±0.08 for MATISSE, and ±0.02 for GRA4MAT. The MATISSE value matches the estimation given in the first light paper (Lopez et al., 2022, Table 5), whereas the GRA4MAT value is better than the MATISSE standalone performance in good atmospheric conditions. Additional information on the raw contrast performance is given in Table 4.
@@ 4.5 Visibility measurements
Absolute visibility measurements with MATISSE require accurate photometry, and hence chopping at frequencies higher than for targets fainter than at and at 4 µm on ATs (respectively and with UTs). This cannot be improved by GRA4MAT. Only the DITs for the low spectral resolution of MATISSE are compatible with this chopping frequency. However, absolute visibility measurements in the continuum are mandatory only in the same low spectral resolution. At higher spectral resolutions, MATISSE users are mostly looking for differential measures between spectral features that are not affected by broad-band photometric errors. Hence, absolute visibility measurements with GRA4MAT should be restricted to low resolution, while higher resolutions with long DITs and without chopping should be used only for coherence flux and differential measurements. Simultaneous chopping and fringe tracking were developed for GRA4MAT in this context (see section 2.6).
@@ 5 Outlook
The performance of GRAVITY for MATISSE presented here corresponds to what was achieved at the end of the project in mid-2023, and justify the recommendation to always use GRA4MAT over MATISSE standalone whenever possible. In this last section, we describe the first science observations carried out with the narrow off-axis mode, as it illustrates nicely the newly acquired capabilities of MATISSE. We then describe how upgrades to the VLTIUT infrastructure are going to further improve GRA4MAT within the next two years. Finally, we suggest additional improvements that are currently not in the VLTI plan but that would further increase the science relevance of MATISSE.
@@ 5.1 Science highlight: A first exoplanet
Soon after the first exoplanet discovery (Mayor & Queloz, 1995), their observation by differential phase interferometry666Other interferometric techniques such as indirect detection by wide angle astrometry (Shao & Colavita, 1992; Colavita et al., 1998, 1999; Delplancke, 2008; Woillez et al., 2010) and direct detection by nulling (Serabyn et al., 2012; Absil et al., 2006) have also been proposed and explored, but were not developed here. was considered (Akeson & Swain, 1999; Petrov et al., 2001). Detection attempts based on the differential phase technique were carried out at VLTI, with the near-infrared AMBER instrument (Millour et al., 2006; Vannier et al., 2006). In the mid-infrared specifically, a first attempt was made with simultaneous MIDI and AMBER observations (Matter et al., 2010). The goals were to evaluate and calibrate the chromatic effects of air, strongly impacting the mid-infrared domain, and to prepare the approach for future MATISSE observations in this field. In addition, it was realised at that time that parasitic light effects would introduce biases affecting the planet signal measurement (Matter et al., 2009); MATISSE was designed to prevent such parasitic effects. The scientific requirements and conceptual study of MATISSE were taking into account the exigent needs imposed by high-contrast measurements on exoplanets (Lopez et al., 2014; Wolf et al., 2016).
@@
The high sensitivity potential of the off-axis fringe tracking with GRA4MAT was demonstrated on the planet $β$ Pictoris b with MATISSE during the commissioning of the GRA4MAT narrow off-axis mode, on the nights of November 8, 2022, and February 3, 2023 (Prog. ID 60.A-9257). These observations were motivated by an earlier observing Guaranteed Observing Time proposal (‘Retrieving the $β$ Pictoris b spectrum in L band’, Prog. ID 110.23SJ), which was submitted to the ESO Observing Program Committee in September 2022, but was not accepted because the mode had not yet been implemented.
We selected the host star $β$ Pictoris a as the fringe-tracking guide star, and alternatively observed the host star and the planet, at the offset predicted by whereistheplanet
@@
Figure 9 shows the fringes obtained on the planet position where the contributions of the L= planet and of the star are nearly equivalent.
This corresponds to a 1000-fold attenuation factor for this L= host star at ( at ).
A preliminary data reduction derived from the K-band GRAVITY approach of GRAVITY Collaboration et al. (2020) yielded a first estimate of the planet spectrum.
To check consistency with a detection of $β$ Pictoris b, we used species
@@
We then compared our preliminary L-band spectrum to this best fit, extrapolated over the L band.
@@ 5.2 Planned improvements
The performance of GRA4MAT will continue to improve due to imminent changes to the VLTI infrastructure for the UTs. The new GRAVITY+ Adaptive Optics (GPAO) for the VLTI/UT, developed by the GRAVITY+ project (GRAVITY+ Collaboration et al., 2022a), will significantly enhance the behaviour of GRA4MAT on the large telescopes. Expected in 2024, the natural guide star (NGS) mode will deliver significantly higher Strehl on brighter objects and make the instrument much less sensitive to degraded seeing conditions. We expect the fringe tracker operation to become fringe-jump-free on the UTs, just as it currently is on the ATs. When the laser guide star (LGS) mode of GPAO becomes available, good Strehl conditions will be obtained on targets that are much fainter in the visible. The fringe tracking limiting magnitude mismatch between ATs and UTs will then be remedied. At a foreseen limiting magnitude of K=, and virtually no limit on the adaptive optics (AO) tip-tilt star at GAIA G=, GRA4MAT will be able to systematically observe red and faint targets, such as young stellar objects and active galactic nuclei.
We also expect that the higher and more stable Strehl delivered to the fringe tracker will make GRA4MAT compatible with chopping on the UTs. This capability will enable the high-precision visibility measurements, in medium resolution and above, required for sensitive mid-infrared interferometric imaging with the larger telescopes.
As explained in section 2.3, GRA4MAT uses the main VLTI delay lines as common-path piston actuators, and is affected by from a latency penalty (see Fig. 3). This penalty will soon be removed by using the beam compressor differential delay lines recently installed by the GRAVITY+ project. These additional common-path delay lines, primarily developed as improved differential delay lines for GRAVITY wide off-axis observations (GRAVITY+ Collaboration et al., 2022b), provide the same level of performance as the TTP internal actuators of the GRAVITY fringe tracker.
@@ 5.3 Possible future developments
At the time of writing this paper, with the exception of section 5.2, the GRA4MAT project at ESO was concluded. What follows are possible developments that should be considered.
@@
The GRA4MAT implementation could benefit from an additional fast piston actuator not in the common path, but exclusive for MATISSE. With such a set-up, the GRAVITY fringe tracker could send an open-loop feed-forward command, improving further the OPD rejection at low frequencies. This control strategy was prototyped (Lane & Colavita, 2003) on the Palomar Testbed Interferometer (Colavita et al., 1999) and operated with the Keck Nuller (Colavita et al., 2010) on the Keck Interferometer (Colavita et al., 2013). The benefit for MATISSE, a higher and more stable fringe contrast, might not be as strong as for a high-precision nulling instrument, for example NOTT (Defrère et al., 2018).
GRA4MAT would also benefit from further developing the MATISSE pipeline. Coherent integrations in post-processing over durations well beyond the DITs are already possible in the N band. In the LM bands, however, longer than coherent integrations must be implemented, especially for observations involving the more sensitive UTs. The $β$ Pictoris b results shown in Fig. 9 actually demonstrate that integrations beyond the DITs are possible. Based on Fig. 6, the instantaneous coherent flux S/N for this target is , whereas the contrast spectrum derived from the combination of many DITs reaches a S/N of . We hope that one day MATISSE becomes capable of measurements on targets virtually invisible in individual LM-band DITs.
We believe that adding a wide off-axis capability to MATISSE has a strong scientific potential. A MATISSE wide implementation would not require any new hardware developments. The dual-field building blocks are already present at VLTI and actively used by GRAVITY wide (GRAVITY+ Collaboration et al., 2022b). MATISSE and the infrared tip-tilt sensor IRIS can already receive the off-axis beams while keeping GRAVITY on axis. MATISSE wide would just be an extension of the GRA4MAT control software and the commissioning of a new mode. Considering that the atmospheric turbulence retains coherence at larger off-axis distances in the mid-infrared, MATISSE wide would drastically expand the MATISSE sky coverage. We would then be able to start observation campaigns of emblematic AGN and YSO science cases (see e.g. Boskri et al., 2021).
@@ 6 Conclusions
The GRAVITY fringe tracker is now used by both the GRAVITY and MATISSE instruments. As a result, MATISSE saved costs and benefited from the fringe tracking expertise developed by the GRAVITY team. Likewise, GRAVITY benefited from a single focus on its fringe tracker, avoiding dispersion and effort duplication. For example, MATISSE motivated the fringe jump control feature, which is also an added value for GRAVITY for high-precision differential visibility across the K band. Therefore, selecting one shared fringe tracker federated the teams around one single project, and made it better for all. This illustrates the value of a mutualisation strategy.
Thanks to GRA4MAT, MATISSE now delivers its predicted performance. A summary is given in Table 4 of all the performance improvements, and of the performance improvements expected with the UTs, when GPAO becomes available in late 2024. A second wave of astrophysical MATISSE observations involving sensitivity, spectroscopy, accuracy, and contrast are now in progress; the associated publications are on their way. The first narrow off-axis observations with MATISSE are opening a new field of mid-infrared extrasolar planet observations with VLTI (Houllé et al., 2024).
For the past 20 years, single-telescope instrumentation for high angular resolution and high-contrast observations has been unimaginable without the support of an AO system (Wizinowich et al., 2000; Brandner et al., 2002). A similar revolution is now happening for optical long-baseline interferometry. This paradigm shift reveals itself clearly when comparing early interferometry reviews (Quirrenbach, 2001; Monnier, 2003) to more recent ones (Eisenhauer et al., 2023). The interferometric instruments, that used to focus on bright targets within the tight coherence limits of the atmosphere, cannot be conceived any more without fringe tracking support. Initiated by the GRAVITY instrument (GRAVITY Collaboration et al., 2017), GRAVITY for MATISSE has now completed this revolution for all modern VLTI instruments.
@@ Acknowledgements.
Based on observations collected at the European Southern Observatory under ESO commissioning programme 60.A-9257. This research has made use of the Jean-Marie Mariotti Center Aspro service.888Available at http://www.jmmc.fr/aspro P. Berio, M. Houllé, S. Lacour and F. Millour acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-21-CE31-0017 (project ExoVLTI). J. Leftley and R. Petrov acknowledge the support of the french ANR under the grant AGN_MELBa ANR ANR-21-CE31-001. A. Meilland is supported by the ANR grant MASSIF (ANR-21-CE31-0018). J. Varga is funded from the Hungarian NKFIH OTKA project no. K-132406. J. Varga has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 716155 (SACCRED). The research of J. Varga is supported by NOVA, the Netherlands Research School for Astronomy. J. Varga acknowledges support from the Fizeau exchange visitors program. The research leading to these results has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement 101004719 (ORP).References